# Video Understanding with Large Language Models: A Survey

### v7

@@ IV-B Open-ended Evaluation
Open-ended evaluation involves questions without pre-defined options or structured formats. While ground-truth answers serve as references, scoring methods are more sophisticated than option selection or string matching. GPT-3.5/4 models often evaluate predictions by comparing them with reference answers. Notable open-ended benchmarks include MovieChat-1K [241], MLVU [294], NExT-QA [111], VELOCITI [305], and EAGLE [306]. These require more complex responses demonstrating deeper reasoning. CinePile [299] incorporates analytical tasks like character dynamics and narrative analysis. The most popular GPT-based evaluation methods, proposed in [99], are Open-end Zero-shot Video QA Evaluation and Video-based Generative Performance Benchmarking. Performance comparisons of Vid-LLMs on these metrics are shown in Table V. Originally closed-ended benchmarks like MSRVTT-QA [93], MSVD-QA [93], TGIF-QA [107], and ActivityNet-QA [110] can be repurposed as open-ended in GPT-based evaluations, as LLMs generate free-form responses that GPT models compare to references. These methods have limitations: evaluation scores change with GPT version updates, making cross-study comparisons difficult; results depend heavily on prompt engineering; and LLM evaluators may favor responses similar to their generation patterns rather than objectively assessing quality.
@@ IV-C Other Evaluations
Other benchmarks evaluate fine-grained temporal and spatiotemporal understanding. Dense captioning generates [122, 123, 124, 125, 126] detailed descriptions for multiple video events/objects, using BLEU, METEOR, and CIDEr metrics that assess both temporal localization and descriptive accuracy. Vid-LLMs’ performance of dense video captioning on ActivityNet Captions [127] is shown in Table IV. Several Vid-LLMs have already achieved performance comparable to traditional task-specific models in dense video captioning. Video temporal grounding localizes specific moments based on textual queries, evaluated using tIoU and Recall@K. Spatiotemporal grounding extends this to localize in both space and time, assessed via spatiotemporal IoU and mAP. Object tracking [251, 256] follows objects across frames, evaluated using precision, success rate, and tracking accuracy. Video saliency detection [307] identifies visually salient regions, evaluated with AUC-J, NSS, etc. These tasks rely on temporal or spatiotemporal annotations as ground-truth, with metrics like IoU, Recall@K, and mAP widely adopted. Human evaluation is also used for subjective aspects, though this is labor-intensive and time-consuming.
As for qualitative evaluation, several approaches can effectively assess Vid-LLMs’ performance in addition to numerical metrics. Error analysis [203, 308] for open-ended QA and difference comparisons [258, 120] between model outputs and ground truth annotations (e.g., intervals) for temporal/spatiotemporal understanding provide insights into model limitations. Attention visualization [309] reveals what visual elements the models prioritize when generating responses. Self-explanation [308, 293], where models justify their answers for closed-ended benchmarks, offers valuable insights into reasoning processes and potential misconceptions. Human studies, though resource-intensive, remain helpful in finding models that reflect human preferences.
@@
Momentor [262] 14.9 2.3 4.7
TimeChat [257] 19.0 4.7 5.7 36.9
VTG-LLM [259] 20.7 5.1 5.9 34.8
AVicuna [120] 22.5 5.1 6.5 45.2
TRACE [312] 25.9 6.0 6.4 39.3
VTimeLLM [313] 27.6 5.8 6.8
TRACE-uni [312] 29.2 6.4 6.9 40.4
GIT [314] 29.8 5.7 7.8 50.6
Vid2Seq [271] 30.2 5.9 8.5 51.8
Streaming Vid2Seq [315] 37.8 6.2 10.0 52.9
Streaming GIT [315] 41.2 6.6 9.0 50.9
@@
GPT4-V [316] - - 59.5 4.09 3.88 4.37 3.94 4.02 4.06
Video-LLaMA [245] 51.6 29.6 12.4 1.96 2.18 2.16 1.82 1.79 1.98
LLaMA-Adapter [317] 54.9 43.8 34.2 2.03 2.32 2.30 1.98 2.15 2.16
VideoChat [270] 56.3 45.0 26.5 2.23 2.50 2.53 1.94 2.24 2.29
VALLY [244] 60.5 51.1 45.1 2.43 2.13 2.86 2.04 2.45 2.38
MovieLLM [223] 63.2 52.1 43.3 2.64 2.61 2.92 2.03 2.43 2.53
Video-ChatGPT [99] 64.9 49.3 35.2 2.50 2.57 2.69 2.16 2.20 2.42
Chat-UniVi [237] 65.0 54.6 46.1 2.89 2.91 3.46 2.89 2.81 2.99
Vista-LLaMA [236] 65.3 60.5 48.3 2.44 2.64 3.18 2.26 2.31 2.57
AV-LLM [231] 67.3 53.7 47.2 2.56 2.47 2.93 2.17 2.47 2.52
LLaVA-NeXT-Video [318] 67.8 - 53.5 3.39 3.29 3.92 2.60 3.12 3.26
LLaMA-VID [238] 69.7 57.7 47.4 2.96 3.00 3.53 2.46 2.51 2.89
VTimeLLM [258] 69.8 58.8 45.5 2.78 3.10 3.40 2.49 2.47 2.85
VideoChat2 [203] 70.0 54.1 49.1 3.02 2.88 3.51 2.66 2.81 2.98
LongVLM [213] 70.0 59.8 47.6 2.76 2.86 3.34 2.39 3.11 2.89
AVicuna [120] 70.2 59.7 53.0 2.81 2.62 3.25 2.53 2.59 2.76
Video-LLaVA [239] 70.7 59.2 45.3 2.87 2.94 3.44 2.45 2.51 2.84
RED-VILLM [211] 71.2 53.9 44.2 2.71 2.75 3.34 2.34 2.45 2.72
Video LLaMA 2 [201] 71.7 - 49.9 3.09 3.09 3.68 2.63 3.25 3.15
Artemis [191] 72.1 56.7 39.3 2.69 2.55 3.04 2.24 2.70 2.64
VideoGPT+ [200] 72.4 60.6 50.6 2.46 2.73 2.81 1.78 2.47 2.45
MiniGPT4-video [215] 73.9 58.8 44.4 3.09 3.02 3.57 2.56 2.67 2.98
ST-LLM [218] 74.6 63.2 50.9 3.23 3.05 3.74 2.93 2.81 3.15
MovieChat [241] 75.2 52.7 45.7 2.76 2.93 3.01 2.24 2.42 2.67
PLLaVA [217] 76.6 62.0 56.3 3.21 2.86 3.62 2.33 2.93 2.99
IG-VLM [159] 76.7 62.7 57.3 3.26 2.76 3.57 2.34 3.28 3.04
@@ V-A1 Media and Entertainment
  • Online Video Platforms and Multimedia Information Retrieval: Vid-LLMs significantly enhance search algorithms [319], generate context-aware video recommendations [320], and aid in natural language tasks such as subtitle generation and translation [271], contributing to online video platforms and multimedia retrieval systems. Their capabilities in analyzing videos for specific keyword retrieval [168, 321, 322] improve intelligent recommendation systems. In the multimedia fields, it combines videos in domains like music [323], avatar [324, 325, 326, 327], and scene [328], to assist with content generation.
  • Video Summarization and Editing: Vid-LLMs are integral in generating concise summaries of video content [329], which analyzes visual and auditory elements to extract key features for context-aware summaries. They also contribute to the field of video editing, as covered in existing literature [75] and advertisement editing [114].
@@ V-A2 Interactive and User-Centric Systems
  • Virtual Education, Accessibility, and Sign Language: Vid-LLMs serve as virtual tutors in education, analyzing instructional videos for interactive learning environments [330]. They also facilitate sign language translation into spoken language or text [331, 332], improving accessibility for the deaf and hard of hearing.
  • Interactive Gaming and Virtual Environments: In the gaming industry, Vid-LLMs play a crucial role in creating dynamic dialogues and storylines, as well as aiding in generating procedural content, such as quests and in-game texts [333, 334]. They also power customer service chatbots [335, 336]. Additionally, in AR/VR/XR, Vid-LLMs contribute to the generation of dynamic narrative content, enhancing user immersion [337, 338, 339, 340].
  • State-Aware Human-Computer Interaction and Robot Planning: In the field of human-computer interaction, Vid-LLMs analyze user videos to discern context and provide customized assistance, as highlighted in Bi et al. [341]. Interaction forms also involve video content understanding like captioning videos [342, 343, 155]. Concurrently, in autonomous robot navigation, the SayPlan method [344] integrates LLMs with 3D scene graphs to enable robots to interpret and navigate complex spaces in large buildings.
@@ V-A3 Healthcare and Security Applications
  • Healthcare Innovations: In the healthcare sector, Vid-LLMs play a crucial role in processing and interpreting medical literature, assisting in diagnostic and educational processes [345, 346, 347, 348], and providing decision support for healthcare professionals. They are used in patient interaction tools, such as chatbots for symptom assessment and addressing health-related queries, thus improving patient care and access to information [349].
  • Security, Surveillance, and Cybersecurity: Vid-LLMs are crucial in security and protection, analyzing communications for potential threats [350, 351] and detecting anomalous patterns in data [352, 353]. In surveillance video analysis, they identify suspicious behaviors, helping law enforcement [354]. Their role in cybersecurity includes identifying phishing attempts and contributing to forensic analysis by summarizing case-related texts [355]. They may also improve video crowd counting [356] for security applications.
  • Autonomous Vehicles: In autonomous vehicles, Vid-LLMs can process language inputs for interaction [357], assist in understanding road signs and instructions [247, 358], and improve user interfaces for vehicle control systems [357], enhancing safety and user experience.
@@ V-A4 Other Applications
Vid-LLMs offer valuable applications beyond those previously discussed. In video generation research [359, 360, 361], Vid-LLMs can evaluate model performance, refine text prompts, and provide reasoning capabilities that better reflect human intentions. Additionally, Vid-LLMs show promise in resource-constrained environments through edge computing applications [362, 363, 364] and can enhance privacy-preserving distributed systems through federated learning frameworks [365, 366, 367].
@@ V-B1 More Fine-grained Video Understanding
Fine-grained understanding remains challenging due to limited datasets, insufficient research, and high computational demands. The frame-by-frame analysis increases computational load while capturing spatiotemporal information. Understanding deeper semantics (emotions, scene dynamics) is harder, though text-video alignment through LLMs offers promise [293].
@@ V-B2 Long-form Video Understanding
Long videos’ extended duration complicates the analysis, especially in understanding events over time. Thus, identifying key events and maintaining attention in long videos is difficult [196, 213, 186]. Effective mechanisms are needed to detect and highlight important parts, particularly in content-rich or complex plot videos.
@@ V-B3 Multimodal Video Understanding
Multimodal video understanding requires integrating different types of data, such as visual, audio, and text, for a better understanding of videos [120, 368, 369]. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. The field lacks research and datasets, with challenges in ensuring high-quality data annotation.
@@ V-B4 Hallucination in Video LLMs
Hallucination occurs when models generate responses disconnected from source material [370], caused by insufficient feature extraction, influence of video context, domain gap between vision and language, and inherent LLM hallucinations. Solutions include post-training strategies [210], enhanced spatiotemporal context understanding, and visual-linguistic latent collaboration.
@@ V-B5 Industrial Deployment and Scalability
Effective deployment strategies [213, 371, 372, 212, 373, 363] include model compression, token merging, domain-specific fine-tuning, modular architectures, efficient caching, and standardized integration frameworks, balancing efficiency with performance for industrial systems.

### v8

@@ IV-B Open-ended Evaluation
Open-ended evaluation involves questions without pre-defined options or structured formats. While ground-truth answers serve as references, scoring methods are more sophisticated than option selection or string matching. GPT-3.5/4 models often evaluate predictions by comparing them with reference answers. Notable open-ended benchmarks include MovieChat-1K [song2023moviechat], MLVU [zhou2024mlvu], NExT-QA [NExT-QA], VELOCITI [saravanan2024velociti], and EAGLE [bi2024eagle]. These require more complex responses demonstrating deeper reasoning. CinePile [rawal2024cinepile] incorporates analytical tasks like character dynamics and narrative analysis. The most popular GPT-based evaluation methods, proposed in [maaz2023video], are Open-end Zero-shot Video QA Evaluation and Video-based Generative Performance Benchmarking. Performance comparisons of Vid-LLMs on these metrics are shown in Table V. Originally closed-ended benchmarks like MSRVTT-QA [xu2017video], MSVD-QA [xu2017video], TGIF-QA [jang2017tgif], and ActivityNet-QA [yu2019activitynet] can be repurposed as open-ended in GPT-based evaluations, as LLMs generate free-form responses that GPT models compare to references. These methods have limitations: evaluation scores change with GPT version updates, making cross-study comparisons difficult; results depend heavily on prompt engineering; and LLM evaluators may favor responses similar to their generation patterns rather than objectively assessing quality.
@@ IV-C Other Evaluations
Other benchmarks evaluate fine-grained temporal and spatiotemporal understanding. Dense captioning generates [shao2022region, Wang_2021_ICCV, shao2023textual, long2023capdet, shao2024dcmstrd] detailed descriptions for multiple video events/objects, using BLEU, METEOR, and CIDEr metrics that assess both temporal localization and descriptive accuracy. Vid-LLMs’ performance of dense video captioning on ActivityNet Captions [krishna2017dense] is shown in Table IV. Several Vid-LLMs have already achieved performance comparable to traditional task-specific models in dense video captioning. Video temporal grounding localizes specific moments based on textual queries, evaluated using tIoU and Recall@K. Spatiotemporal grounding extends this to localize in both space and time, assessed via spatiotemporal IoU and mAP. Object tracking [wang2024elysium, Li2024GroundingGPT] follows objects across frames, evaluated using precision, success rate, and tracking accuracy. Video saliency detection [tang2024cardiff] identifies visually salient regions, evaluated with AUC-J, NSS, etc. These tasks rely on temporal or spatiotemporal annotations as ground-truth, with metrics like IoU, Recall@K, and mAP widely adopted. Human evaluation is also used for subjective aspects, though this is labor-intensive and time-consuming.
As for qualitative evaluation, several approaches can effectively assess Vid-LLMs’ performance in addition to numerical metrics. Error analysis [li2024mvbench, hua2024mmcomposition] for open-ended QA and difference comparisons [huang2023vtimellm, tang2024avicuna] between model outputs and ground truth annotations (e.g., intervals) for temporal/spatiotemporal understanding provide insights into model limitations. Attention visualization [bi2024unveiling] reveals what visual elements the models prioritize when generating responses. Self-explanation [hua2024mmcomposition, tang2024vidcomposition], where models justify their answers for closed-ended benchmarks, offers valuable insights into reasoning processes and potential misconceptions. Human studies, though resource-intensive, remain helpful in finding models that reflect human preferences.
@@
Momentor [Qian2024Momentor] 14.9 2.3 4.7
TimeChat [Ren2024TimeChat] 19.0 4.7 5.7 36.9
VTG-LLM [guo2024vtgllm] 20.7 5.1 5.9 34.8
AVicuna [tang2024avicuna] 22.5 5.1 6.5 45.2
TRACE [guo2024trace] 25.9 6.0 6.4 39.3
VTimeLLM [huang2024vtimellm] 27.6 5.8 6.8
TRACE-uni [guo2024trace] 29.2 6.4 6.9 40.4
GIT [wang2022git] 29.8 5.7 7.8 50.6
Vid2Seq [yang2023vid2seq] 30.2 5.9 8.5 51.8
Streaming Vid2Seq [zhou2024streamingdvc] 37.8 6.2 10.0 52.9
Streaming GIT [zhou2024streamingdvc] 41.2 6.6 9.0 50.9
@@
GPT4-V [gpt4v2023online] - - 59.5 4.09 3.88 4.37 3.94 4.02 4.06
Video-LLaMA [zhang2023video] 51.6 29.6 12.4 1.96 2.18 2.16 1.82 1.79 1.98
LLaMA-Adapter [gao2023llama] 54.9 43.8 34.2 2.03 2.32 2.30 1.98 2.15 2.16
VideoChat [li2023videochat] 56.3 45.0 26.5 2.23 2.50 2.53 1.94 2.24 2.29
VALLY [luo2023valley] 60.5 51.1 45.1 2.43 2.13 2.86 2.04 2.45 2.38
MovieLLM [chen2023videollm] 63.2 52.1 43.3 2.64 2.61 2.92 2.03 2.43 2.53
Video-ChatGPT [maaz2023video] 64.9 49.3 35.2 2.50 2.57 2.69 2.16 2.20 2.42
Chat-UniVi [Jin2024Chat-UniVi] 65.0 54.6 46.1 2.89 2.91 3.46 2.89 2.81 2.99
Vista-LLaMA [Ma2023Vista-LLaMA] 65.3 60.5 48.3 2.44 2.64 3.18 2.26 2.31 2.57
AV-LLM [Shu2023AV-LLM] 67.3 53.7 47.2 2.56 2.47 2.93 2.17 2.47 2.52
LLaVA-NeXT-Video [zhang2024llavanextvideo] 67.8 - 53.5 3.39 3.29 3.92 2.60 3.12 3.26
LLaMA-VID [li2023llama-vid] 69.7 57.7 47.4 2.96 3.00 3.53 2.46 2.51 2.89
VTimeLLM [huang2023vtimellm] 69.8 58.8 45.5 2.78 3.10 3.40 2.49 2.47 2.85
VideoChat2 [li2024mvbench] 70.0 54.1 49.1 3.02 2.88 3.51 2.66 2.81 2.98
LongVLM [weng2024longvlm] 70.0 59.8 47.6 2.76 2.86 3.34 2.39 3.11 2.89
AVicuna [tang2024avicuna] 70.2 59.7 53.0 2.81 2.62 3.25 2.53 2.59 2.76
Video-LLaVA [lin2023video] 70.7 59.2 45.3 2.87 2.94 3.44 2.45 2.51 2.84
RED-VILLM [huang2024redvillm] 71.2 53.9 44.2 2.71 2.75 3.34 2.34 2.45 2.72
Video LLaMA 2 [cheng2024videollama2] 71.7 - 49.9 3.09 3.09 3.68 2.63 3.25 3.15
Artemis [qiu2024artemis] 72.1 56.7 39.3 2.69 2.55 3.04 2.24 2.70 2.64
VideoGPT+ [maaz2024videogpt+] 72.4 60.6 50.6 2.46 2.73 2.81 1.78 2.47 2.45
MiniGPT4-video [ataallah2024minigpt4video] 73.9 58.8 44.4 3.09 3.02 3.57 2.56 2.67 2.98
ST-LLM [liu2024stllm] 74.6 63.2 50.9 3.23 3.05 3.74 2.93 2.81 3.15
MovieChat [song2023moviechat] 75.2 52.7 45.7 2.76 2.93 3.01 2.24 2.42 2.67
PLLaVA [xu2024pllava] 76.6 62.0 56.3 3.21 2.86 3.62 2.33 2.93 2.99
IG-VLM [Kim2024IG-VLM] 76.7 62.7 57.3 3.26 2.76 3.57 2.34 3.28 3.04
@@ V-A1 Media and Entertainment
  • Online Video Platforms and Multimedia Information Retrieval: Vid-LLMs significantly enhance search algorithms [mao2023large], generate context-aware video recommendations [ju2022prompting], and aid in natural language tasks such as subtitle generation and translation [yang2023vid2seq], contributing to online video platforms and multimedia retrieval systems. Their capabilities in analyzing videos for specific keyword retrieval [zhao2023lavila, jin2023text, jin2023diffusionret] improve intelligent recommendation systems. In the multimedia fields, it combines videos in domains like music [xu2023launchpadgpt], avatar [song2021tacr, song2023emotional, song2024tri, song2021fsft], and scene [Song_2023_CVPR], to assist with content generation.
  • Video Summarization and Editing: Vid-LLMs are integral in generating concise summaries of video content [Pramanick_2023_ICCV], which analyzes visual and auditory elements to extract key features for context-aware summaries. They also contribute to the field of video editing, as covered in existing literature [wu2023visual] and advertisement editing [tang2022multi].
@@ V-A2 Interactive and User-Centric Systems
  • Virtual Education, Accessibility, and Sign Language: Vid-LLMs serve as virtual tutors in education, analyzing instructional videos for interactive learning environments [gan2023large]. They also facilitate sign language translation into spoken language or text [liu2023survey, de2023machine], improving accessibility for the deaf and hard of hearing.
  • Interactive Gaming and Virtual Environments: In the gaming industry, Vid-LLMs play a crucial role in creating dynamic dialogues and storylines, as well as aiding in generating procedural content, such as quests and in-game texts [mishra2023generating, koomen2023text]. They also power customer service chatbots [soni2023large, medeiros2023analysis]. Additionally, in AR/VR/XR, Vid-LLMs contribute to the generation of dynamic narrative content, enhancing user immersion [gokce2023role, jung2023xr, yu2024promptfix, huang2023egocentric].
  • State-Aware Human-Computer Interaction and Robot Planning: In the field of human-computer interaction, Vid-LLMs analyze user videos to discern context and provide customized assistance, as highlighted in Bi et al. [bi2023misar]. Interaction forms also involve video content understanding like captioning videos [wang2023caption, hu2022promptcap, hua2024finematch]. Concurrently, in autonomous robot navigation, the SayPlan method [rana2023sayplan] integrates LLMs with 3D scene graphs to enable robots to interpret and navigate complex spaces in large buildings.
@@ V-A3 Healthcare and Security Applications
  • Healthcare Innovations: In the healthcare sector, Vid-LLMs play a crucial role in processing and interpreting medical literature, assisting in diagnostic and educational processes [eysenbach2023role, liu2022beat, liu2022disco, liu2024emage], and providing decision support for healthcare professionals. They are used in patient interaction tools, such as chatbots for symptom assessment and addressing health-related queries, thus improving patient care and access to information [li2023llava].
  • Security, Surveillance, and Cybersecurity: Vid-LLMs are crucial in security and protection, analyzing communications for potential threats [al2023chatgpt, mouratidis2023modelling] and detecting anomalous patterns in data [lee2023lanobert, almodovar2023logfit]. In surveillance video analysis, they identify suspicious behaviors, helping law enforcement [de2023socratic]. Their role in cybersecurity includes identifying phishing attempts and contributing to forensic analysis by summarizing case-related texts [tang2023graphgpt]. They may also improve video crowd counting [cao2025efficientmaskedautoencodervideo] for security applications.
  • Autonomous Vehicles: In autonomous vehicles, Vid-LLMs can process language inputs for interaction [cui2024drive], assist in understanding road signs and instructions [li2023otter, lai2023lisa], and improve user interfaces for vehicle control systems [cui2024drive], enhancing safety and user experience.
@@ V-A4 Other Applications
Vid-LLMs offer valuable applications beyond those previously discussed. In video generation research [zhou2024survey, lin2023videodirectorgpt, kondratyuk2023videopoet], Vid-LLMs can evaluate model performance, refine text prompts, and provide reasoning capabilities that better reflect human intentions. Additionally, Vid-LLMs show promise in resource-constrained environments through edge computing applications [jin2024efficient, lu2024b, hu2023edge] and can enhance privacy-preserving distributed systems through federated learning frameworks [yao2024federated, bastola2024fedmil, wang2023fedvmr].
@@ V-B1 More Fine-grained Video Understanding
Fine-grained understanding remains challenging due to limited datasets, insufficient research, and high computational demands. The frame-by-frame analysis increases computational load while capturing spatiotemporal information. Understanding deeper semantics (emotions, scene dynamics) is harder, though text-video alignment through LLMs offers promise [tang2024vidcomposition].
@@ V-B2 Long-form Video Understanding
Long videos’ extended duration complicates the analysis, especially in understanding events over time. Thus, identifying key events and maintaining attention in long videos is difficult [zhang2024longva, weng2024longvlm, Wang2024LifelongMemory]. Effective mechanisms are needed to detect and highlight important parts, particularly in content-rich or complex plot videos.
@@ V-B3 Multimodal Video Understanding
Multimodal video understanding requires integrating different types of data, such as visual, audio, and text, for a better understanding of videos [tang2024avicuna, mohammadkhani2025survey, zhangmultimedia]. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. The field lacks research and datasets, with challenges in ensuring high-quality data annotation.
@@ V-B4 Hallucination in Video LLMs
Hallucination occurs when models generate responses disconnected from source material [zhang2024eventhallusion], caused by insufficient feature extraction, influence of video context, domain gap between vision and language, and inherent LLM hallucinations. Solutions include post-training strategies [zhang2024llavahounddpo], enhanced spatiotemporal context understanding, and visual-linguistic latent collaboration.
@@ V-B5 Industrial Deployment and Scalability
Effective deployment strategies [weng2024longvlm, lee2024video, tang2024enhancing, tan2024koala, shang2024interpolating, lu2024b] include model compression, token merging, domain-specific fine-tuning, modular architectures, efficient caching, and standardized integration frameworks, balancing efficiency with performance for industrial systems.
@@ VI Conclusion
Yunlong Tang received the B.Eng. degree in Intelligence Science and Technology from the Southern University of Science and Technology (SUSTech) in 2023, supervised by Prof. Feng Zheng. She is currently pursuing a Ph.D. degree in Computer Science at the University of Rochester, advised by Prof. Chenliang Xu. Her research focuses on multimodal learning, especially video understanding and generation.
Jing Bi is currently pursuing a Ph.D. in Computer Science at the University of Rochester since 2020, advised by Prof. Chenliang Xu. He received his B.S. from Shandong University and M.S. from the University of Rochester.
Siting Xu received her B.Eng. (2019 - 2023) degree in Computer Science and Technology from Southern University of Science and Technology (SUSTech), supervised by Prof. Feng Zheng.
Luchuan Song is currently a Ph.D. candidate in Computer Science at the University of Rochester. He received his M.S. and B.S. from University of Science and Technology of China.
Susan Liang is currently a Ph.D. candidate in the Computer Science Department at the University of Rochester. His research focuses on multimodal learning.
Teng Wang is currently a Ph.D. candidate in the Department of Computer Science at the University of Hong Kong. He obtained his bachelor’s and master’s degrees from Sun Yat-sen University in 2017 and 2020, respectively. His research interests lie in vision-language multimodal learning and video understanding.
Daoan Zhang is currently a Ph.D. Student in Computer Science at the University of Rochester, advised by Prof. Jiebo Luo. His research focuses on generative AI.
Jie An is a Ph.D. candidate in Computer Science at the University of Rochester, advised by Prof. Jiebo Luo. Prior to that, he obtained his B.S. (2012 - 2016) and M.S. (2016 - 2019) in Applied Mathematics from Peking University, advised by Prof. Jinwen Ma. Jie’s research focuses on improving the performance and extending the capability of GenAI models. He is particularly interested in image style transfer, generative models, image/video generation, and multi-modal generation/evaluation.
Jingyang Lin is a PhD student majoring in Computer Science at the University of Rochester, advised by Professor Jiebo Luo. He received his BE and MSc degrees from Sun Yat-sen University (SYSU), Guangzhou, China, in 2019 and 2022, respectively. His research interests include multimodal learning with LLMs, AI for health, and self-supervised learning.
Rongyi Zhu is currently a Ph.D. student in Computer Science at Stony Brook University. He received his MS. degree in Computer Science from the University of Rochester in 2024. His research focuses on trustworthy AI.
Ali Vosoughi is a PhD candidate in Electrical and Computer Engineering at the University of Rochester, working with Professors Chenliang Xu and Axel Wismueller. His research focuses on using AI for multimodality and complex reasoning to assist humans with challenging tasks. He holds a Bachelor’s degree in Electrical Engineering from Sharif University of Technology (Iran), and two Master’s degrees—one from Bogazici University (Turkey) and another from the University of Rochester (USA).
Chao Huang is currently a PhD candidate in Computer Science at the University of Rochester, advised by Prof. Chenliang Xu. Previously, he spent a year as a research assistant at the Chinese University of Hong Kong, working with Prof. Chi-Wing Fu. He received his B.Eng. from ESE Department, Nanjing University in 2019.
Zeliang Zhang received a B.Eng. degree in computer science from Huazhong University of Science and Technology in 2022. He is currently a Ph.D. candidate in Computer Science at the University of Rochester. His research focuses on trustworthy and efficient deep learning methods.
Pinxin Liu is a Ph.D. student at the University of Rochester, advised by Prof. Chenliang Xu. Before that, he was a Research Scientist Intern at Flawless AI. He received his B.S. from the Department of Computer Science at the University of Rochester. His research interest lies in human-related topics, e.g., video gesture synthesis, 3D face rendering, and text-to-motion generation.
Mingqian Feng received his B.S. degree in Physics from University of Science and Technology of China (USTC) and his M.S.E. degree in financial mathematics from Johns Hopkins University. He is currently a Ph.D. student in Computer Science at the University of Rochester. His research focuses on bias, optimization, and video understanding.
Feng Zheng is an Associate Professor at Southern University of Science and Technology (SUSTech). His research interests include machine learning (ML), computer vision (CV) and human-computer interaction (HCI). He received a Ph.D. from the University of Sheffield, UK. Before joining SUSTech, he worked as a senior researcher at Tencent YouTu Lab in Shanghai, China. Prior to this, he worked as a postdoctoral researcher at the University of Pittsburgh, USA and as an assistant research professor at the Shenzhen Institute of Advanced Technology, CAS. In terms of academic research, he has published 85 papers in top international journals and conferences, including IEEE TPAMI/TITS/TIP, AAAI, NeuIPS, CVPR, and ECCV.
Jianguo Zhang is a Professor in the Department of Computer Science and Engineering, Southern University of Science and Technology. Previously, he was a Reader in Computing, School of Science and Engineering, University of Dundee, UK. He received a PhD from the National Lab of Pattern Recognition, Institute of Automation, Chinese Academy of Sciences, Beijing, China, 2002. His research interests include object recognition, medical image analysis, machine learning, and computer vision. He is a senior member of the IEEE and an Associate Editor of IEEE Transactions on Multimedia.
Ping Luo is an Associate Professor in the Department of Computer Science at the University of Hong Kong, Associate Director of the HKU Musketeers Foundation Institute of Data Science, and Deputy Director of the Joint Research Lab of HKU and Shanghai AI Lab. He earned his Ph.D. in Information Engineering from the Chinese University of Hong Kong in 2014, supervised by Prof. Xiaoou Tang and Prof. Xiaogang Wang. Before joining HKU in 2019, he was a Research Director at SenseTime. He has published over 100 papers in top conferences and journals, with 50,000+ citations on Google Scholar. His awards include the 2015 AAAI Easily Accessible Paper, 2022 ACL Outstanding Paper, 2023 WAIC Outstanding Paper, and ICCV’23 Best Paper nomination. In 2020, he was named one of the MIT Technology Review’s Innovators Under 35 in Asia-Pacific. He has mentored 30 Ph.D. students, many of whom have won prestigious awards like the Nvidia and Baidu Fellowships.
Jiebo Luo is the Albert Arendt Hopeman Professor of Engineering at the University of Rochester, which he joined in 2011 after over 15 years at Kodak Research Laboratories. He has played key roles in numerous conferences, serving as General Co-Chair of ACM Multimedia 2018 and IEEE ICME 2024, and Program Co-Chair of ACM Multimedia 2010, IEEE CVPR 2012, and IEEE ICIP 2017. He has served on editorial boards of several major journals, including IEEE Transactions on Pattern Analysis and Machine Intelligence and IEEE Transactions on Multimedia, and was the Editor-in-Chief of IEEE Transactions on Multimedia from 2020 to 2022. Prof. Luo is a Fellow of ACM, AAAI, IEEE, AIMBE, IAPR, and SPIE, and a member of Academia Europaea (AE) and the US National Academy of Inventors (NAI). His honors include the 2024 IEEE Computer Society Edward J. McClusky Technical Achievement Award, the 2021 ACM SIGMM Technical Achievement Award, the 2024 William H. Riker University Award for Excellence in Graduate Teaching, the 2024 Edmund A. Hajim Outstanding Faculty Award, and the inaugural 2024 Debra Haring Excellence in Research Award.
Chenliang Xu is an Associate Professor in the Department of Computer Science at the University of Rochester. He received his Ph.D. in Computer Science from the University of Michigan in 2016, an M.S. in Computer Science from the University at Buffalo in 2012, and a B.S. in Information and Computing Science from Nanjing University of Aeronautics and Astronautics, China, in 2010. His research originates in computer vision and tackles interdisciplinary topics, including video understanding, audio-visual learning, vision and language, and methods for trustworthy AI. Xu is a recipient of the James P. Wilmot Distinguished Professorship (2021), the University of Rochester Research Award (2021), the Best Paper Award at the 17th ACM SIGGRAPH VRCAI Conference (2019), the Best Paper Award at the 14th Sound and Music Computing Conference (2017), and the University of Rochester AR/VR Pilot Award (2017). He has authored over 100 peer-reviewed papers in computer vision, machine learning, multimedia, and AI venues. He served as an associate editor for IEEE Transactions on Multimedia and area chair/reviewer for various international conferences.